# Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities

### v2

@@ 5.1 Complementing Foundation Models with Practical Alternatives
Based on our observations from our corpus of research papers, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs). We group existing approaches into several broad categories, which allows us to track adoption trends and the rise of PLMs, beginning with the use of BERT-based models in finance around 2019 (Araci, 2019a; Yang et al., 2020c), and accelerating rapidly with the emergence of general-purpose LLMs like ChatGPT (Wu et al., 2023; Gao et al., 2023). However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems. These factors make non-PLM methods (such as rule-based classifiers, feature-engineered models, and domain-specific embeddings) not only viable but often preferable in production environments (Shah et al., 2023; Lopez et al., 2021). The rest of this section revisits key methodological alternatives to PLMs, each addressing practical and conceptual needs that remain critical in industrial financial tasks.
@@
Classical NLP pipelines are far more resource-efficient than LLMs. In time-time contexts like news analytics or alert systems, lightweight models (e.g., logistic regression or small CNNs) deliver insights instantly and can scale to massive corpora without requiring GPU infrastructure (Zhai et al., 2019). Systems like RavenPack (Cornez, 2018) and Refinitiv News Analytics (TRNA) (Waeosri, 2020) use rule-based NLP and horizontally scalable architectures to deliver structured news data with sub-second latency, supporting real-time trading and high-throughput analysis.
@@
Lexicon- and rule-based methods remain essential in regulated settings. Financial sentiment lexicons (e.g., Loughran-McDonald) (Loughran and McDonald, 2011) and keyword rules are transparent, auditable, and often outperform neural models in compliance or risk-flagging tasks (Hosseini et al., 2018). Lexicon-based surveillance tools, such as those used in communications monitoring (NICE Actimize, 2023) and financial crime compliance (Risk & Compliance, 2021), are widely adopted for their transparency and ease of audit, allowing firms to justify alerts with specific keywords.
@@
Named-entity and relation extraction systems power knowledge graphs and schema-driven applications. These systems offer precision, structure, and explainability, which are ideal for fraud detection, Know Your Customer (KYC), and regulatory alignment, where PLMs often lack controllability (Szarvas et al., 2007; Tarnopolski et al., 2019). Financial institutions use rule-based NLP for entity and relation extraction to build and update knowledge graphs in KYC and Anti-Money Laundering (AML) workflows, as seen in solutions from KYCHub (2025) and TigerGraph (2024), enabling automated risk monitoring and explainable link discovery.
@@
Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited data. They enable explicit domain knowledge integration (such as tagging financial terms or numerical cues) for credit scoring, bankruptcy prediction, and event detection (Li et al., 2018; Khandani et al., 2021; Alanis et al., 2022). Financial firms often use TF-IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
@@
Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents. These methods are fast, interpretable, and competitive on factual corpora like earnings calls or filings (Blei et al., 2003; Nenkova and McKeown, 2011). Companies like BlackRock, Bank of America (Phoenix Strategy Group, 2025), and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
@@ 5.2 Implications of Shift Toward General-Purpose Models
General-purpose language models have rapidly become the default choice in NLP applied to Finance, with models like RoBERTa (Liu et al., 2019), GPT-4 (Achiam et al., 2023), and LLaMA-2 (Touvron et al., 2023) adopted far more widely and quickly than domain-specific alternatives. As Figure 6 shows, even earlier versions of FinBERT (Araci, 2019b; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022) are used more than newer ones, and recent finance-tuned models like FinGPT (Yang et al., 2023) remain rare. This trend reflects how easy it has become to apply strong general models out of the box. For practitioners, this means many models are optimized for benchmarks, not for the realities of finance, where interpretability, latency, or regulation often matter more than raw accuracy.
@@
This shift is also mirrored in the decline of custom architectures. Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020b)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)). Today, these efforts are rarely pursued, which seems like an efficiency gain at first glance. However, as Section 5.1 discusses, real-world financial applications often require design trade-offs, such as transparency for compliance, or fast inference for deployment, and off-the-shelf LLMs do not accommodate them well. The move away from custom models narrows the solution space at a time when financial tasks still demand methodological flexibility.
@@ 5.3 Enabling Reproducibility through Open Resources
As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets.
@@ 6 Conclusion & Discussion
Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance-specific evaluation metrics, stress-testing on crisis data, and truly global, multilingual datasets. These findings underscore both the field’s progress and the work ahead to build resilient, transparent, and inclusive financial NLP solutions. Below, we discuss their implications for researchers and practitioners.
@@
The application of LMs to financial domains presents challenges that differ from traditional NLP benchmarks. Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant, even amid the LLM wave. The lack of financial evaluation metrics, underuse of crisis-period data, and limited attention to multilingual and global financial corpora signal areas where NLP researchers can contribute meaningfully. Moreover, the scarcity of domain-adapted models points to missed opportunities for more effective domain transfer. Future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.

### v3

@@ 5.1 Complementing Foundation Models with Practical Alternatives
Based on our observations from our vast corpus of research, the methodological landscape in NLP for Finance has shifted decisively toward pretrained language models (PLMs) and large language models (LLMs). We group existing approaches into several broad categories, which allows us to track adoption trends and the rise of PLMs, beginning with the use of BERT-based models in finance around 2019 (Araci, 2019a; Yang et al., 2020c), and accelerating rapidly with the emergence of general-purpose LLMs like ChatGPT (Wu et al., 2023; Gao et al., 2023). However, NLP tasks in Finance often pose domain-specific constraints such as interpretability, regulatory transparency, data scarcity, and the need for low-latency systems Luo et al. (2018); Maia et al. (2018). These factors make non-PLM methods (such as rule-based classifiers, feature-engineered models, and domain-specific embeddings) not only viable but often preferable in production environments (Shah et al., 2023; Lopez et al., 2021). The rest of this section revisits key methodological alternatives to PLMs, each addressing practical and conceptual needs that remain critical in industrial financial tasks.
@@
Classical NLP pipelines are far more resource-efficient than LLMs Strubell et al. (2019). In time-time contexts like news analytics or alert systems, lightweight models (e.g., logistic regression or small CNNs) deliver insights instantly and can scale to massive corpora without requiring GPU infrastructure (Zhai et al., 2019). Systems like RavenPack (Cornez, 2018) and Refinitiv News Analytics (TRNA) (Waeosri, 2020) use rule-based NLP and horizontally scalable architectures to deliver structured news data with sub-second latency, supporting real-time trading and high-throughput analysis.
@@
Lexicon- and rule-based methods remain essential in regulated settings Moreno-Ortiz et al. (2020); Du et al. (2023). Financial sentiment lexicons (e.g., Loughran-McDonald) (Loughran and McDonald, 2011) and keyword rules are transparent, auditable, and often outperform neural models in compliance or risk-flagging tasks (Hosseini et al., 2018). Lexicon-based surveillance tools, such as those used in communications monitoring (NICE Actimize, 2023) and financial crime compliance (Risk & Compliance, 2021), are widely adopted for their transparency and ease of audit, allowing firms to justify alerts with specific keywords.
@@
Named-entity and relation extraction systems power knowledge graphs and schema-driven applications Elhammadi et al. (2020); Hamad et al. (2024). These systems offer precision, structure, and explainability, which are ideal for fraud detection, Know Your Customer (KYC), and regulatory alignment, where PLMs often lack controllability (Szarvas et al., 2007; Tarnopolski et al., 2019). Financial institutions use rule-based NLP for entity and relation extraction to build and update knowledge graphs in KYC and Anti-Money Laundering (AML) workflows, as seen in solutions from KYCHub (2025) and TigerGraph (2024), enabling automated risk monitoring and explainable link discovery.
@@
Feature-based pipelines using TF-IDF, static embeddings, and classical classifiers (SVMs, XGBoost) work well with limited dataWang and Manning (2012). They enable explicit domain knowledge integration (such as tagging financial terms or numerical cues) for credit scoring, bankruptcy prediction, and event detection (Li et al., 2018; Khandani et al., 2021; Alanis et al., 2022). Financial firms often use TF‑IDF (Simha, 2021) and XGBoost (Kamau, 2024) to build effective models with limited labeled data without requiring costly infrastructure.
@@
Unsupervised techniques like LDA and extractive summarization remain valuable for discovering latent topics or summarizing dense financial documents Filippova et al. (2009); Agrawal et al. (2021). These methods are fast, interpretable, and competitive on factual corpora like earnings calls or filings (Blei et al., 2003; Nenkova and McKeown, 2011). Companies like Bank of America (Phoenix Strategy Group, 2025) and American Express (Arora and Radhakrishnan, 2020) apply topic modeling and extractive summarization to earnings calls and filings, using unsupervised or minimally supervised methods for reliable and traceable outputs.
@@ 5.2 Implications of Shift Toward General-Purpose Models
General-purpose language models have rapidly become the default choice in NLP applied to Finance, with models like RoBERTa (Liu et al., 2019), GPT-4 (Achiam et al., 2023), and LLaMA-2 (Touvron et al., 2023) adopted far more widely and quickly than domain-specific alternatives. As Figure 7 shows, even earlier versions of FinBERT (Araci, 2019b; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022) are used more than newer ones, and recent finance-tuned models like FinGPT (Yang et al., 2023) remain rare. For practitioners, this means many models are optimized for benchmarks, not for the realities of finance, where interpretability, latency, or regulation often matter more than raw accuracy.
@@
This shift is also mirrored in the decline of custom architectures. Prior to 2022, task-specific models were more common: TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020c)); MDRM (Qin and Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021) in volatility prediction (Niu et al., 2023; Mathur et al., 2022a)). Today, these efforts are rarely pursued. However, as Section 5.1 discusses, real-world financial applications often require design trade-offs, such as transparency for compliance, or fast inference for deployment, and off-the-shelf LLMs do not accommodate them well. The move away from custom models narrows the solution space at a time when financial tasks still demand methodological flexibility.
@@ 5.3 Enabling Reproducibility through Open Resources
As we witness an accelerated adoption of language models, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets.
@@ 6 Conclusion & Discussion
Drawing on Table 3, our survey shows that sentiment analysis, information extraction, and forecasting have driven NLP in finance, yet critical gaps remain: finance‑specific evaluation metrics, stress‑testing on crisis data, and truly global, multilingual datasets. These findings underscore both the field’s progress and the work ahead to build resilient, transparent, and inclusive financial NLP solutions.
@@
The application of LMs to financial domains presents challenges that differ from traditional NLP benchmarks. Our findings suggest that domain-specific requirements, such as interpretability, latency, and compliance, make classical or hybrid approaches highly relevant. The lack of financial evaluation metrics, underuse of crisis-period data, and limited attention to multilingual and global financial corpora signal areas where NLP researchers can contribute meaningfully. Moreover, the scarcity of domain-adapted models points to missed opportunities for more effective domain transfer. As a result, our observations indicate that future work should prioritize robustness, transparency, and real-world relevance over benchmark performance alone.
@@
Limitations
For readability, we standardize task names and merge closely related variants when counting. When a paper spans multiple areas, we assign a single primary tag to keep statistics interpretable; a multi-tag alternative would yield slightly different totals. These are intentional choices to aid comparability.
Ethical Considerations
This study does not assess the merit or quality of individual papers. We do not suggest that any category or method is inherently better than others. Our goal is to map the volume and distribution of research applying NLP methods to finance, and to offer a foundation for further study, without making judgments about the value of specific approaches.
Appendix A Timeline Figure
This appendix includes the timeline Figure 7.
Appendix B Authorship and Funding
Based on the analysis we conducted, we also made a binary classification of the following categories:
  • Academic authors
  • Industrial authors
  • Funding (industrial)
  • Funding (governmental)
  • Funding (academic)
As seen in the Figure 6, we see that there is no major correlation between funding and authorship. Thus it does not imply that if an industrial/academic author is present in the list of authors, there will necessarily be funding from the industry/academia/government. This finding is surprising actually as one would expect that a paper written by only academics would have academic/governmental funding and the same for industrial authors and industrial funding.
Table 4: Summary of Authors and Funding Categories
Category Sum
Academic and Industrial 78
Only Academic authors 118
Only Industrial authors 25
Funding (industrial) 49
Funding (governmental) 61
Funding (academic) 35
Appendix C Survey Dimensions, Methodologies, and Task Categories in Financial NLP Research
Figure 6: This figure displays the correlation between authors from industrial and academic backgrounds within various types of funding such as academic, industrial and governmental.
This appendix consolidates the key reference tables from our study, providing an overview of the dimensions examined (Table 5), the methodologies employed (Table 6), and the task categories identified in NLP research applied to the financial domain (Tables 7 and 8).
Table 5: This table outlines the dimensions examined in our analysis of NLP research focused on the financial domain. Dimensions that led to significant findings are discussed in the main paper, while an analysis of authorship and funding is provided in Appendix B.
Exploration dimension Description
Primary task The main task addressed in the paper, such as volatility prediction (Wang et al., 2024a).
Sub-Tasks Additional tasks that support the primary objective, like using sentiment analysis to enhance stock price predictions (Jain and Agrawal, 2024).
Methodology Techniques applied, ranging from traditional machine learning to deep learning and large language models. See Section LABEL:sec:methodologies.
Code and data availability Whether the paper provides open-source code or data, and the quality of this accessibility (e.g., active links).
Contribution What type of contribution does the paper have (Dataset, framework, model, evaluation, etc.)?
Comparability of research What other research did the researchers compare their work to?
Data source The types of data used, including financial reports, social media, news articles, or other financial documents.
Metrics Evaluation criteria, including standard ML metrics (accuracy, precision) and finance-specific metrics (Sharpe ratio).
Data years The time periods covered by the data, such as crisis years or more stable financial periods.
Authorship Indicates if the paper was authored or co-authored by industry professionals, reflecting practical applications.
Funding source The origin of the research funding (academic, industrial, or governmental) and its influence on the research focus.
Table 6: This table categorizes the key methodologies into nine groups.“PLMs” stands for Pretrained Language Models, while “LLMs” stands for Large Language Models.
Methodology Description
Statistical NLP Methods like TF-IDF, Bag-of-Words, and n-grams focused on extracting statistical patterns from financial text.
Embeddings Techniques such as Word2Vec, GloVe, and custom embeddings designed to map financial terms into vector spaces, improving downstream tasks.
Conventional ML Algorithms like Logistic Regression, Support Vector Machines (SVM), and Decision Trees, often used for classification and risk prediction.
Deep Learning (DL) Custom neural network architectures tailored for specific financial tasks such as stock price forecasting.
PLMs Encoder-only models (e.g., BERT) and Encoder-Decoder models (e.g., T5) are used for tasks like document summarization and information extraction.
LLMs Models like GPT-3.5 and LLaMA-2 focus on text generation and understanding complex financial language.
Statistical Modeling Involves correlation analysis, Granger causality, etc., to understand relationships in financial data.
Graphs Uses graph structures to model interactions in financial systems.
Knowledge Graphs Integrating structured knowledge within financial tasks to enhance model performance and reliability.
Primary Task Examples & Insights
Finance-Related (Categories IV)
Explainable AI (XAI) refers to methods that make machine learning outputs more interpretable and less of a "black box." These papers show how small input changes can drastically affect model explanations (Sinha et al., 2021) and highlight the instability of tools like LIME in sensitive applications (Burger et al., 2023). In finance, where decisions must be traceable and justifiable, clear and reliable explanation methods are essential for building trust in tasks like fraud detection and risk assessment.
Privacy-preserving methods protect sensitive data while still allowing model training and analysis. Examples include homomorphic encryption for text similarity (Kim et al., 2022a), federated learning for distributed training without raw data exchange (Zhao et al., 2024c), and domain adaptation that shares only model parameters (Xiao et al., 2024). These methods are especially important in finance, where institutions must work with private data while staying compliant and avoiding breaches.
Numerical reasoning involves working with numbers and structured data to solve problems or develop algorithms. This includes solving math word problems using declarative knowledge (Roy and Roth, 2018), learning numeracy through number embeddings (Duan et al., 2021), and pretraining models for verifying tabular claims (Akhtar et al., 2023). In finance, good numerical reasoning supports tasks like market prediction and risk evaluation by improving how models interpret and reason about numbers.
Financial Applications (Category III)
Sentiment analysis helps assess market mood and forecast financial movements using sources like news, social media, and reports. Studies analyze opinions in reports and news (Rodriguez Inserte et al., 2023), investor sentiment on social media (Guo et al., 2023), and tone in financial texts (Choe et al., 2023). As markets become more influenced by public opinion, sentiment analysis plays a growing role in trading and analysis.
Information extraction (IE) and relation extraction focus on identifying entities and linking them to events or other entities. Examples include extracting financial events from documents (Zheng et al., 2019), signals in reports (Huang et al., 2023), and patterns in social media (Conforti et al., 2022). These tools help connect companies, individuals, and financial instruments to real-world developments (Liou et al., 2021), supporting real-time analysis and prediction.
Question Answering (QA) enables systems to find specific information in financial documents, reports, and databases. QA models support quick access to relevant facts (Mavi et al., 2023), and recent work focuses on calibrating large language models (LLMs) for finance-specific tasks (Zhao et al., 2024b; Theuma and Shareghi, 2024; Addlesee, 2024). This area emphasizes not only building QA tools but also improving their accuracy for use in high-stakes financial contexts.
Table 7: Overview of key primary tasks in Categories IV (Finance-Related) above the mid line and Category III (Financial Applications) below. Each entry summarizes the task’s role and relevance in NLP research applied to finance, with representative examples and practical insights.
Primary Task Examples & Insights
Financial Resources (Category II)
Dataset and Resource Construction involves building structured, labeled financial datasets for model development and evaluation. These papers focus on creating large-scale annotated resources (Chen et al., 2020a), including datasets like Tab-CQA (Liu et al., 2023) and ConvFinQA (Chen et al., 2022b) designed for reasoning over tables and multi-step numerical queries. Other works assemble corpora from financial reports (Shah et al., 2022), company filings (Zmandar et al., 2022), news (Tang et al., 2023), and government documents (Shah et al., 2023), enabling downstream tasks like entity linking, numerical reasoning, or forecasting.
Fraud detection leverages domain-specific data and models to identify deceptive financial behavior across platforms. Erben and Waldis (2024) identifies financial scams on Instagram using a fine-tuned BERT model deployed via browser extension and REST API. Another approach detects identity fraud through interactive dialogue, employing knowledge graphs and reinforcement learning to expose inconsistencies in claimed personal data (Wang et al., 2019). These systems highlight how NLP architectures can protect users from financial manipulation.
Financial Forecast (Category I)
Stock Price and Volatility Prediction aims to forecast stock movements or market instability using text data. These models use historical data (Sawhney et al., 2021b), news (Ahbali et al., 2022), and event-driven signals (Sawhney et al., 2020a; Wu, 2020) to predict stock trends. Volatility prediction further explores market sensitivity by analyzing unstructured text like press releases or financial articles (Qin and Yang, 2019), helping traders anticipate fluctuations.
Risk Prediction involves identifying the likelihood and impact of adverse financial events, such as defaults or market disruptions. These models analyze unstructured data like earnings call transcripts (Sang and Bao, 2022), regulatory filings, and legal documents to detect early signals of financial risk (Li et al., 2023a). Applications include credit risk estimation, fraud detection, and systemic risk monitoring (Zhang et al., 2024).
Table 8: Descriptions of major primary tasks in Category II (Financial Resources) above the line and Category I (Financial Forecast) below. Each entry summarizes the task’s focus and contribution to NLP research applied to finance, with representative examples.
Appendix D Conference Proceedings
Table 9 categorizes key conferences and workshops according to the four primary task categories identified in our survey.
Appendix E Papers per year
Table 10 reports the number of papers per year from 1975 to 2024.
Appendix F Glossary
Table 11 provides definitions to the key financial metrics mentioned in this paper.
Category Conferences and Workshops
Financial Forecast (Category I) ACL, Australasian Language Technology Association Workshop, CCL, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), Financial Technology and Natural Language Processing (FinNLP) Workshop, ICCL, LREC-COLING, NAACL, NAACL (Industry), SRW
Financial Resources (Category II) ACL, ACL (Industry), CoNLL, EACL, EcoNLP Workshop, EMNLP, EMNLP-IJCNLP, e-Commerce and NLP Workshop, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, GWC, ICON, ICCL, LREC, LREC-COLING, NAACL, NAACL (Industry), NLP4PI Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, SIGHUM Workshop
Financial Applications (Category III) ACL, ATALA, Bridging Human–Computer Interaction and Natural Language Processing Workshop, CCL, CoLM, Computational Approaches to Subjectivity, Sentiment and Social Media Analysis Workshop, DeeLIO Workshop, EACL, EcoNLP Workshop, EMNLP, EMNLP (Industry), EMNLP-IJCNLP, Financial Narrative Processing and MultiLing Financial Summarisation Workshop, Financial Technology and Natural Language Processing (FinNLP) Workshop, ICON, ICCL, INLG, LREC, LREC-COLING, NAACL, NAACL (Industry), Natural Legal Language Processing Workshop, News Media Content Analysis and Automated Report Generation Workshop, Pattern-based Approaches to NLP in the Age of Deep Learning Workshop, Safety4ConvAI Workshop, Structured Prediction for Natural Language Processing Workshop, TextGraphs
Finance Related (Category IV) ACL, ACL-IJCNLP, BlackboxNLP Workshop, EMNLP, Financial Technology and Natural Language Processing (FinNLP) Workshop, LREC-COLING, SRW, TrustNLP Workshop
Table 9: Conference and Workshop Categorization
Year 2024 2023 2022 2021 2020 2019 2018 2017 2016 2015 2014 2012 2010 2008 2006 2002 2001 1976 1975
Number of Papers 127 53 90 32 57 15 14 21 6 1 5 1 3 3 1 1 1 1 1
Table 10: Number of papers satisfying our filtration criterion per year.
Metric Definition
Sharpe Ratio Measures how much return an investment gives for each unit of risk, compared to a risk-free asset (Investopedia, 2025a).
Maximum Drawdown The biggest drop from a peak to a low point in a portfolio’s value before it recovers.
Cumulative Return The total profit or loss from an investment over time.
Table 11: Key Definitions of financial terms used in our paper.
Appendix G Data Sources
Document type Sources Accessibility (incl. access method) Modalities Update cadence
Regulatory filings SEC EDGAR; other registries: SEDAR (Canada), EDINET (Japan), Companies House (UK) Public via web UI; free EDGAR APIs and RSS (U.S. SEC, 2024); other registries public but less commonly used Text/PDF Event-driven; periodic (annual/quarterly)
Earnings calls (transcripts, slides, audio) Seeking Alpha (Seeking Alpha, 2025); The Motley Fool (The Motley Fool, 2025); FactSet (FactSet Insight, 2025); Thomson Reuters (Thomson Reuters, 2025); S&P Global (S&P Global Market Intelligence, 2025); company IR sites Mixed: many providers behind paywalls; IR sites often public; access via web UI and provider APIs (where offered) Text transcripts; audio Quarterly around earnings; event-driven (scheduled call dates)
Financial news Reuters, Bloomberg, Dow Jones, CNBC, FT, WSJ; free aggregators like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025) Mixed: premium (e.g., Bloomberg) and free sources, including static datasets (UC Irvine, 2023; Zhou et al., 2024); access via web UIs and vendor APIs (e.g., Bloomberg API (Bloomberg L.P., 2025a)) Text (articles; headlines) Real-time / continuous
Social media X (Twitter); Reddit; sentiment datasets for X (Kaggle Contributor, 2025; Chaudhary, 2020; Xu and Cohen, 2018b) and Reddit (Kaggle Contributor, 2021; Wang et al., 2024b) X API with limited free and paid tiers (X (2025), formerly Twitter); StockTwits free/commercial API (StockTwits, 2025); public web UIs Short-text posts; sentiment tags (StockTwits Bullish/Bearish) Real-time
Macroeconomic & monetary communications U.S. Federal Reserve (FOMC) statements and transcripts (Federal Reserve System, 2025) Public via official web portals; archives available Text (policy statements, minutes, press releases) Scheduled (policy meetings, release calendars)
Economic indicators BEA (U.S. BEA, 2025), FRED (Federal Reserve Bank of St. Louis, 2024), the World Bank (The World Bank, 2025) Public via official portals and APIs Numeric time series + accompanying text commentary Scheduled (monthly / quarterly)
Analyst reports Bloomberg Terminal (Bloomberg L.P., 2025b); S&P Capital IQ (S&P Global, 2025) Paid/proprietary subscription platforms; enterprise access/APIs; occasional public excerpts in news/Seeking Alpha Text/PDF (reports, notes) Ongoing; event-driven (coverage initiations, earnings, sector updates)
Table 12: Financial text/data sources by document type with sources, accessibility (and access method), modalities, and update cadence.
Regulatory Filings and Earnings Calls
U.S. public companies file a range of documents such as 10‑K, 10‑Q, 8‑K, credit agreements, proxy statements, S‑1, S‑3, and others are accessible via the SEC’s EDGAR platform, through free API calls and RSS feeds (U.S. SEC, 2024). While similar registries exist (e.g., Canada’s SEDAR, Japan’s EDINET, UK’s Companies House), they are rarely used, reinforcing a U.S.-centric and English-language bias (Section 4.4). Earnings calls, where firms discuss quarterly performance with investors, are another rich data source Mukherjee et al. (2022); Koval et al. (2023); Pardawala et al. (2025). Seeking Alpha transcribes  4,500 calls quarterly (Seeking Alpha, 2025); other providers include The Motley Fool (The Motley Fool, 2025), FactSet (FactSet Insight, 2025), Thomson Reuters (Thomson Reuters, 2025), and S&P Global (S&P Global Market Intelligence, 2025), though often behind paywalls. Companies also post transcripts/slides on IR websites. Public datasets include transcriptions and audio, such as S&P 500 calls from 2017 (Qin and Yang, 2019) and 1,213 companies’ data covering 2015-2018 (Li et al., 2020b).
Financial News and Social Media
Real-time news from sources like Reuters, Bloomberg, Dow Jones, CNBC, FT, and WSJ is vital for forecasting. The Reuters-21578 dataset (UC Irvine, 2023), with 10,369 articles from 1987, remains a benchmark in text analysis. Larger archives such as FNSPID (Zhou et al., 2024) link 15.7 million articles (1999-2023) to S&P 500 stocks. News is accessible via APIs (e.g., Bloomberg (Bloomberg L.P., 2025a)) and free sources like Yahoo Finance (Yahoo Finance, 2025) and Google News (Google News, 2025). On social media, “FinTwit” and platforms like X and StockTwits are used for sentiment analysis. Key datasets include Sentiment140 (Kaggle Contributor, 2025; Chaudhary, 2020), StockNet (Xu and Cohen, 2018b), and Reddit datasets (Kaggle Contributor, 2021; Wang et al., 2024b). X’s API offers limited free and full historical access under paid tiers (X (2025), formerly Twitter). StockTwits, with Bullish/Bearish sentiment tagging, offers free and commercial API access (StockTwits, 2025) and is often used in forecasting studies.
Macroeconomic and Monetary Data
Central bank communications (policy statements, minutes, and press releases) offer critical macroeconomic guidance Ahrens and McMahon (2021); Peskoff et al. (2023); Menzio et al. (2024b). The U.S. Federal Reserve’s FOMC releases detailed statements and transcripts on interest rates, inflation, and economic outlook (Federal Reserve System, 2025), with archives available on the Fed’s website. Despite their market relevance, such texts remain underused in NLP-based forecasting. Shah et al. (2023) addresses this by introducing an annotated dataset of FOMC communications and proposing a sentence-level classification task (hawkish, dovish, neutral) using fine-tuned transformers. Economic indicators like GDP, CPI, and unemployment figures – issued by the BEA (U.S. BEA, 2025), FRED (Federal Reserve Bank of St. Louis, 2024), Eurostat (Eurostat, 2025), IMF (International Monetary Fund, 2025), OECD (Organisation for Economic Co-operation and Development, 2025), and the World Bank (The World Bank, 2025) – are mostly numerical but often include press commentary, all freely available via official portals.
Analyst Reports